An electron moving at 4.15 x 103 m/s in a 1.05 T magnetic field experiences a magnetic force of 1.35 x 10-16N. 


What angle (in degrees, between 0 and 90°) does the velocity of the electron make with the magnetic field? 
Answer #1

here,

the speed of electron , v = 4.15 * 10^3 m/s

the magnetic feild , B = 1.05 T

the magnetic force , F = 1.35 * 10^-16 N

the charge on electron , q = 1.6 * 10^-19 C

let the angle between velocity of electron and magnetic feild be theta

F = q * v * B * sin(theta)

1.35 * 10^-16 = 1.6 * 10^-19 * 4.15 * 10^3 * 1.05 * sin(theta)

solving for theta

theta = 11.16 degree

the angle between velocity of electron and magnetic feild is 11.16 degree